Shift report 20010815 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Aug 15 2001 - 01:18:40 EDT

  • Next message: Apache: "Shift report 20010815 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Aug 15 01:18:39 2001, Shift: 16:00-24:00
    Supervisor: sjs, with: sjs,ah,hi
    Log book#7 pages 89 - 89
    
    Shift Summary:
    ------------------------------------------------------------
    No beam for most of shift due to DX power supply problem.
    Cogged beam achieved late in shift.
    
    Set low-gain ZDC pedestal locations to ch 10-30.  Also
    attempted to set high-gain ZDC pedestals, but unable to
    locate setting that placed these within adc range. Checked
    Q-L suppression feature on ADC (which is enabled), but
    changing this setting did not seem to have a noticable
    effect.  Decided this exercise better done without beam.
    
    No new detector problems found. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Aug 15 2001 - 01:19:43 EDT